# Privacy-Preserving UCB Decision Process Verification via zk-SNARKs

### v1

@@ 1 Introduction
Research Challenges. Thus, to integrate the UCB algorithm with zk-SNARKs to ensure the data and parameters’ confidentiality while supporting secure UCB decision process verification, we need to follow the above conversion process to convert the UCB algorithm into a format compatible with zero-knowledge proofs, thus facilitating efficient verification. However, this integration presents three major challenges: (1) The inherent randomness of the UCB complicates this task; the challenge is how to transform this randomness into the deterministic process necessary for constructing the arithmetic circuit; (2) Given that zk-SNARK relies on polynomials in finite fields Operations, how to convert the non-polynomial operations of the UCB (such as logarithms and non-integer powers) into polynomial operations is the second challenge. (3) To meet the finite field requirements of R1CS, how to adjust the floating point numbers in the UCB to meet the finite field requirements of R1CS is the third challenge.
@@ 1.2 Use Case
Figure. 2 illustrates the application of zk-SNARKs to the AI doctor using the UCB algorithm. Here, the AI doctor functions as the prover, while the verifier could be either the patient themselves or a third party, such as an insurance company. Initially, the AI doctor gathers the patient’s information, denoted as $x$, and computes a diagnostic result $y$ using the $zkUCB$ model and parameters $a$. Subsequently, the AI doctor creates a common reference string $crs$ and formulates commitments $C_{x}$ and $C_{a}$, based on $x$ and $a$ respectively. The next step involves the AI doctor generating a statement $\phi=(C_{x},C_{a},y)$ and a witness $w=(x,a)$, with the latter serving as a confidential input in the subsequent phase. The AI doctor then produces a proof $\pi=(crs,\phi,w)$ and dispatches $\pi$, the $crs$, and the statement $\phi$ to the third party for verification $VRF=(crs,\phi,\pi)$. After verification, the results $VRF$ are relayed to the patient. This process ensures that commitments to the parameters and private data, along with the computational results and generated proofs, are presented to third parties for verification, all while safeguarding the integrity and confidentiality of both the data and the algorithm.
@@ 1.3 Organization
The subsequent sections of this paper are structured as follows: In Section 2, we provide the essential preliminaries. Section 3 delves into an in-depth discussion of zkUCB. Moving forward to Section 4, we present a thorough examination of comprehensive experimental results and analysis. In Section 5, we review the related work, highlighting the distinctions between our approach and existing methodologies. Finally, Section 6 encapsulates the paper with concluding remarks.
@@ 2.2 A General zk-SNARKs Scheme
Zero-Knowledge Succinct Non-Interactive Argument of Knowledge (zk-SNARKs) Chen et al. (2022) is a novel cryptographic protocol that allows a prover to authenticate a statement’s truth to a verifier without extra information disclosure. zk-SNARKs stand out for their non-interactive nature, eliminating the need for continuous communication typical in traditional proofs. Their succinct proofs, small and fast to verify, remain efficient regardless of the complexity of the statement being proven, enhancing their practicality in diverse applications.
@@
    Setup Phase: the zk-SNARKs scheme employs the setup algorithm to generate a common reference string ($crs$) including the evaluation key and verification key, which is used for proof generation and proof verification.
@@ 3.1.1 Pseudo-random Number Generator
In the field of stochastic processes, the generation of discrete random variables conforming to a uniform distribution can be achieved by inputting a seed parameter into a linear congruential generator (LCG) denoted as $F$. The seed is a parameter determined by the user. In the execution of $F$, there is also a second input parameter $u$, which is the upper bound of its output. Consequently, the sequence of random numbers $\{X_{1},...,X_{n}\}$ with length $n$ is generated by $X_{i}=F(seed,u),1\leq i\leq n$. Every $X_{i}$ belongs to the uniform distribution with mean $u$ and range of $(0,2u)$.
@@ 3.1.2 Non-polynomial Functions
In the context of zk-SNARK, all programs must be performed by polynomial operations in finite fields. A critical aspect of zkUCB involves converting the non-polynomial operations in the UCB algorithm, such as logarithms and non-integer powers, into polynomial forms. Our solution for handling logarithmic functions is the implementation of piecewise linear approximation by focusing on discrete intervals of positive integers. After performing the approximation, and rounding down the results to the nearest integer, we achieved a practical and statistically robust approximation. Similarly, for approximating non-integer powers, particularly square roots as required in the UCB algorithm, we utilized Newton’s method, fixed the number of iterations to $20$, and rounded down the results to the nearest integer. These strategies allow us to approximate these functions using polynomial methods, ensuring compatibility with the computational constraints of zk-SNARKs while maintaining uniform time and space complexity.
@@ 4.3.3 Components Size in zkUCB
Figure 6: The average time during various phases
@@ 5 Related Work
Verifiable Computation: verifiable computation is a key research area essential for validating the correctness of outsourced computations, particularly in cloud services. It focuses on verifying third-party computations without needing to re-execute them fully. This field encompasses a blend of cryptographic and non-cryptographic approaches to ensure computational integrity and data authenticity. Cryptographic methods include Zero-Knowledge Proofs Yang and Wang (2022); Mouris and Tsoutsos (2021); Wahby et al. (2018), Interactive Proofs (IP) Baelde et al. (2021), and Homomorphic Encryption (HE) Kim et al. (2023), which offer robust security guarantees. In contrast, non-cryptographic methods, notably Authenticated Data Structures (ADS) Sun et al. (2020), concentrate on maintaining data integrity. Three primary streams dominate this field: ADS, which ensures data authenticity; IP, which facilitates interactive validation; and zk-SNARKs Groth (2016); Wahby et al. (2018), renowned for efficient, non-interactive proof generation. These diverse methods highlight the field’s complexity and richness, underpinning its significance in areas ranging from cloud computing to privacy-preserving protocols. This comprehensive approach to verifying computational processes without complete re-execution sets the foundation for advanced research and practical applications in computational integrity.
Verifiable Machine Learning: the integration of verifiable computation with machine learning (ML) stands as a transformative advancement, poised to redefine data security and integrity in AI-driven applications. This convergence is not only enhancing the trustworthiness of ML models but is also unlocking new avenues for secure and private data analysis, particularly in sensitive sectors like healthcare, finance, and government. Traditional ML models often operate as black boxes, with limited ability to verify the internal processes and outputs. The advent of verifiable machine learning aims to bring transparency and reliability into this domain. Homomorphic encryption (HE) Li et al. (2020) and garbled circuits (GC) Jawurek et al. (2013) are prevalent methods in this sphere, enabling secure computation of encrypted data. While these methods offer robust privacy guarantees, they often come with computational overheads. Another trend is the use of Zero-Knowledge Proofs in ML Weng et al. (2021); Xing et al. (2023), which allows the validation of model predictions without revealing the underlying data or model parameters. This is particularly useful in scenarios where data confidentiality is paramount. Current research integrating ZKP has primarily concentrated on specific algorithms, including CNN Liu et al. (2021); Lee et al. (2024), DNN Ghodsi et al. (2017), linear regression, logistic regression, neural networks, support vector machines, K-means Zhao et al. (2021), and decision trees Zhang et al. (2020).
@@ 6 Conclusion
This study presents a significant advancement in the integration of reinforcement learning with data privacy through the development of the zkUCB algorithm. By integrating zk-SNARKs within the UCB algorithm, zkUCB ensures the confidentiality and security of sensitive data and parameters, while facilitating a transparent decision-making process. Our comprehensive experimental analysis confirms the superiority of zkUCB over the standard UCB algorithm, particularly in terms of reward optimization and operational efficiency. This improvement is because appropriate quantization bits can effectively reduce information entropy in the decision-making process, enhancing rewards. In addition, the scalability and practicality of zkUCB are demonstrated through its linear proof size and verification times, underlining its potential applicability in a wide range of real-world scenarios. This study thus contributes a pivotal solution to the challenge of balancing data confidentiality with algorithmic transparency and efficiency.

### v3

@@ 1 Introduction
Research Challenges.
Thus, to integrate the UCB algorithm with zk-SNARKs to ensure the data and parameters’ confidentiality while supporting secure UCB decision process verification, we need to follow the above conversion process to convert the UCB algorithm into a format compatible with zero-knowledge proofs, thus facilitating efficient verification. However, this integration presents three major challenges: (1) The inherent randomness of the UCB complicates this task; the challenge is how to transform this randomness into the deterministic process necessary for constructing the arithmetic circuit; (2) Given that zk-SNARK relies on polynomials in finite fields Operations, how to convert the non-polynomial operations of the UCB (such as logarithms and non-integer powers) into polynomial operations is the second challenge. (3) To meet the finite field requirements of R1CS, how to adjust the floating point numbers in the UCB to meet the finite field requirements of R1CS is the third challenge.
@@ 1.2 Use Case
Figure. 2 illustrates the application of zk-SNARKs to the AI doctor using the UCB algorithm. Here, the AI doctor functions as the prover, while the verifier could be either the patient themselves or a third party, such as an insurance company. Initially, the AI doctor gathers the patient’s information, denoted as $x$, and computes a diagnostic result $y$ using the $zkUCB$ model and parameters $a$. Subsequently, the AI doctor creates a common reference string $crs$ and formulates commitments $C_{x}$ and $C_{a}$, based on $x$ and $a$ respectively. The next step involves the AI doctor generating a statement $\phi=(C_{x},C_{a},y)$ and a witness $w=(x,a)$, with the latter serving as a confidential input in the subsequent phase. The AI doctor then produces a proof $\pi=(crs,\phi,w)$ and dispatches $\pi$, the $crs$, and the statement $\phi$ to the third party for verification $VRF=(crs,\phi,\pi)$. After verification, the results $VRF$ are relayed to the patient. This ensures that commitments to the parameters and private data, along with the computational results and generated proofs, are presented to third parties for verification, all while safeguarding the integrity and confidentiality of both the data and algorithm.
@@ 1.3 Organization
The subsequent sections of this paper are structured as follows: In Section 2, we provide essential preliminaries. Section 3 delves into an in-depth discussion of zkUCB. Moving forward to Section 4, we present a thorough examination of comprehensive experimental results and analysis. In Section 5, we review the related work, highlighting the distinctions between our approach and existing methodologies. Finally, Section 6 encapsulates the paper with concluding remarks.
@@ 2.2 A General zk-SNARKs Scheme
Zero-Knowledge Succinct Non-Interactive Argument of Knowledge (zk-SNARKs) Chen et al. (2022) is a novel cryptographic protocol allowing a prover to authenticate a statement’s truth to a verifier without extra information disclosure. zk-SNARKs stand out for their non-interactive nature, eliminating the need for continuous communication typical in traditional proofs. Their succinct proofs, small and fast to verify, remain efficient regardless of the complexity of the proven statement, enhancing their practicality in diverse applications.
@@
    Setup Phase: zk-SNARKs scheme employs the setup algorithm to generate a common reference string ($crs$) including the evaluation key and verification key, which is used for proof generation and proof verification.
@@ 3.1.1 Pseudo-random Number Generator
In the field of stochastic processes, the generation of discrete random variables conforming to a uniform distribution can be achieved by inputting a seed parameter into a linear congruential generator (LCG) denoted as $F$. The seed is a parameter determined by the user. In the execution of $F$, there is also a second input parameter $u$, which is the upper bound of its output. Consequently, the sequence of random numbers $\{X_{1},...,X_{n}\}$ with length $n$ is generated by $X_{i}=F(seed,u),1\leq i\leq n$. Every $X_{i}$ belongs to the uniform distribution with mean $u/2$ and range of $(0,u)$.
@@ 3.1.2 Non-polynomial Functions
In the context of zk-SNARK, all programs must be performed by polynomial operations in finite fields. A critical aspect of zkUCB involves converting non-polynomial operations in the UCB algorithm, such as logarithms and non-integer powers, into polynomial forms. Our solution for handling logarithmic functions is the implementation of piecewise linear approximation by focusing on discrete intervals of positive integers. After performing the approximation, and rounding down results to the nearest integer, we achieved a practical and statistically robust approximation. Similarly, for approximating non-integer powers, particularly square roots as required in the UCB algorithm, we utilized Newton’s method, fixed the number of iterations to $20$, and rounded down results to the nearest integer. These strategies allow us to approximate these functions using polynomial methods, ensuring compatibility with the computational constraints of zk-SNARKs while maintaining uniform time and space complexity.
@@ 4.3.3 Components Size in zkUCB
Figure 6: The size of various components in zkUCB
@@ 5 Related Work
Verifiable Computation.
Verifiable computation is a key research area essential for validating the correctness of outsourced computations, particularly in cloud services. It focuses on verifying third-party computations without needing to re-execute them fully. This field encompasses a blend of cryptographic and non-cryptographic approaches to ensure computational integrity and data authenticity. Cryptographic methods include Zero-Knowledge Proofs Yang and Wang (2022); Mouris and Tsoutsos (2021); Wahby et al. (2018), Interactive Proofs (IP) Baelde et al. (2021), and Homomorphic Encryption (HE) Kim et al. (2023), which offer robust security guarantees. In contrast, non-cryptographic methods, notably Authenticated Data Structures (ADS) Sun et al. (2020), concentrate on maintaining data integrity. Three primary streams dominate this field: ADS, which ensures data authenticity; IP, which facilitates interactive validation; and zk-SNARKs Groth (2016); Wahby et al. (2018), renowned for efficient, non-interactive proof generation. These diverse methods highlight the field’s complexity and richness, underpinning its significance in areas ranging from cloud computing to privacy-preserving protocols. This comprehensive approach to verifying computational processes without complete re-execution sets the foundation for advanced research and practical applications in computational integrity.
Verifiable Machine Learning.
The integration of verifiable computation with machine learning (ML) stands as a transformative advancement, poised to redefine data security and integrity in AI-driven applications. This convergence is not only enhancing the trustworthiness of ML models but is also unlocking new avenues for secure and private data analysis, particularly in sensitive sectors like healthcare, finance, and government. Traditional ML models often operate as black boxes, with limited ability to verify the internal processes and outputs. The advent of verifiable machine learning aims to bring transparency and reliability into this domain. Homomorphic encryption (HE) Li et al. (2020) and garbled circuits (GC) Jawurek et al. (2013) are prevalent methods in this sphere, enabling secure computation of encrypted data. While these methods offer robust privacy guarantees, they often come with computational overheads. Another trend is the use of Zero-Knowledge Proofs in ML Weng et al. (2021); Xing et al. (2023), which allows the validation of model predictions without revealing the underlying data or model parameters. This is particularly useful in scenarios where data confidentiality is paramount. Current research integrating ZKP has primarily concentrated on specific algorithms, including CNN Liu et al. (2021); Lee et al. (2024), DNN Ghodsi et al. (2017), linear regression, logistic regression, neural networks, support vector machines, K-means Zhao et al. (2021), and decision trees Zhang et al. (2020).
@@ 6 Conclusion
This study presents a significant advancement in the integration of reinforcement learning with data privacy through the development of the zkUCB algorithm. By integrating zk-SNARKs within the UCB algorithm, zkUCB ensures the confidentiality and security of sensitive data and parameters, while facilitating a transparent decision-making process. Our comprehensive experimental analysis confirms the superiority of zkUCB over the standard UCB algorithm, particularly in terms of reward optimization and operational efficiency. This improvement is because appropriate quantization bits can effectively reduce information entropy in the decision-making process, enhancing rewards. In addition, the scalability and practicality of zkUCB are demonstrated through its linear proof size and verification times, underlining its potential applicability in a wide range of real-world scenarios.
Acknowledgements
This work is supported in part by the Shanghai Science and Technology Innovation Action Plan 23511100400, in part by the 2023-2024 Open Project of Key Laboratory Ministry of Industry and Information Technology-Blockchain Technology and Data Security 20242216, in part by the Data+ program: TRACK & TRACE – Traceable Pharmaceutical Products project and PAPRiCaS: Programming technology foundations for Accountability, Privacy-by-design & Robustness in Context-aware Systems by Independent Research Fund Denmark.